OpenStudy (anonymous):

find lcm of z^3+8z^2+16z & z^2-8z i got z(z-8)(z+4)^2?

OpenStudy (anonymous):

do i have to put it in decending order?

OpenStudy (anonymous):

z^3+8z^2+16z = z(z^2 + 8z + 16) = z(z+4)^2 z^2 - 8z = z(z-8) lcm = z(z-8)(z+4)^2 - yes

OpenStudy (anonymous):

no

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

yw